Item 6 Part A In Youngs two-slit experiment, the first dark fringe above the central bright fringe occurs at an angle of 0.54 What is the ratio of the slit separation, d to the wavelength of the light、A7 Express your answer using two significant figures 困? Submit My Answers Give Up
Answer #1

We know

  d \sin \theta=(m-\frac{1}{2})\lambda

Given \theta=0.54^o

  m=1

\frac{d}{\lambda}=\frac{(1-\frac{1}{2})}{sin 0.54}

{\color{Red} \frac{d}{\lambda}=53.05} (ans)